help me slove this question . this is from california geomenrty book .

1.3.6.10.15 ...... what is the next pattern ? how did you slove it ?

please help me
thank you in advance

Answer by likaaka(51)

1.3.6.10.15 ......
The answer is 21
You keep adding by the next sequential number
0 + 1 = 1
1 + 2 = 3
3 + 3 = 6
6 + 4 = 10
10 + 5 = 15
15 + 6 = 21
21 + 7 = 28, etc.